OpenStudy (anonymous):

Please walk me through- solve the quadratic equation: 6x=4x^2+1

OpenStudy (anonymous):

Move the 6x to the other side. 4x^2 -6x +1 = 0 Use the Quadratic Formula to find the roots, since you can't factor this easily. x=(-b +/- sqrt(b^2-4ac)) / 2a x=(6 +/- sqrt(36-16)) / 8 = (6+/- sqrt(20)) / 8 = (6 +/- 2*sqrt(5)) / 8 = (3 +/- sqrt(5)) / 4